A Bibliography of Publications of David R. Kincaid

David R. Kincaid University of Texas at Austin Department of Mathematics

Austin, TX USA

Tel: +1 512 471 1242 FAX: +1 512 471 9038

E-mail: kincaid@cs.utexas.edu (Internet)

15 August 2025 Version 1.28

Abstract

This bibliography records publications of David R. Kincaid.

Title word cross-reference

2 [RK92]. GCR(k) [AKMK02]. LU [AKK06, KO88a, OK88]. n [Kin67].

-decomposition [AKK06]. -Factorization [OK88].

10-12 [BS84]. 1983 [BS84, ESY84]. 1984 [Lak84]. 1991 [BD92].

2-4 [BD92]. 2003 [FDKN04]. 205 [KY83]. 2A [KGY80]. 2B [KGRY81]. 2C [KRYG82, KORY84]. 2D [KOY89, KR91, RK91].

539 [LHKK79a]. 586 [KRYG82]. 5th [FDKN04].

Accelerated [KGY80, KGRY81, KRYG82]. Accelerating [YJK84]. Acceleration [Kin81]. Adapting [KY81, KOY82, KY83]. Adaptive [GKMY78, KG77, KGMY79, KGY80, KGRY81, KRYG82]. Advanced [FDKN04, KCO^{+}84]. Afternotes [Kin97]. Algebra [LHKK79a, LHKK79b, BD92]. algorithm [KO88a, KRYG82, LHKK79a]. Algorithmic [Kin67]. Algorithms [GKMY78, KG77, KGMY79, KY81, KY83, KO88b, KO89, OK88, PMKL02, AKMK02, KC01]. alternating [YK95, YK96]. alternating-type [YK95, YK96]. Analysis [Kin71a, Kin71b, KC91b, KC93b, KC93a, Lak84, KC91a, KC96, KC02, KC09, Kin97]. Application [Kin74a]. Applications [FDKN04, RSY93]. applied [AKMK02]. April [BD92]. Architectures [KCO^{+}84]. Arising [OK87c]. Arlington [Lak84]. August [ESY84, FDKN04]. Austin [KH90, KC01].

Basic [LHKK79a, LHKK79b]. Belgium [BD92]. BLAS [OK90, OK94]. Book [Kin94a, Kin97]. Branch [BS84]. brief [KY88]. Brussels [BD92].

C [Kin94a]. California [BS84]. CDC [KY83]. Celebrating [Kin04]. Chebyshev [RSY93]. Class [Kin71a, Kin71b, Kin71c, Kin73b, KY75, Kin73a, YK95, YK96]. Combining [KCO^{+}84]. Common [Kin67]. Comparison [KO85]. Comparisons [KEG^{+}79]. Complex [Kin74b, Kin74a]. Computation [Kin94a]. Computations [KOY84, KOY86a]. Computer [KY75, KOY82, KCO^{+}84, SK91]. Computers [KO85, OK87c, YOKH85, YHOK85]. Computing [CK80a, CK80b, CK85a, CK85b, CK94a, CK94b, Kin67, KC91b, KC93b, KC93a, KC96, KC02, KC09]. Conference [BS84, ESY84, FDKN04, Lak84]. Conferences [KE98]. Conjugate [Kin81, KOJ88, KOJ89b, KOJ89c, OK87a, OK87b, OJK89]. Cray [KO88b, KO89, RK91, RK92]. CYBER [KY83]. Cyclic [SK91].

D [RK92]. David [AK10, Kin04]. decomposition [AKK06]. dedicated [AK10, RSY93]. Degree [Kin74b, YK90, KY91, KY93, Kin94b, Kin71d, Kin74a, KY92, YK72, YK93]. Dense [OK88]. Developed [KY81]. Development [KY75]. Difference [KY86, SK91]. Differential [KY75, KCO^{+}84, KOY86b, YK84b, YK85]. Divisor [Kin67]. Dr. [KVW10].

Efficient [KY83]. Element [KCO^{+}84]. Elliptic [BS84, BS84]. embedded [PMKL02]. ENUMATH [FDKN04]. Equations [Kin71a, Kin71b, Kin71c, Kin71d, Kin73b, KY75, KCO^{+}84, KOY86b, YK84b, Kin73a, YK85]. European [FDKN04]. Experiments [OK87a, OK87b]. Extrapolated [Kin71d].

F1 [LHKK79a]. Factorization

[OK88, KO88a]. Fifty [Kin04]. Finite [KCO^{+}84, KY86, SK91]. Fixed [KY72, YK71]. FORTRAN [KRYG82, KGY80, KGRY81, LHKK79a, LHKK79b]. Four [KO88b, KO89]. Future [KY84, KHY94a].

G. [Kin97]. general [KO88a].

Generalizations [CKY99, LCK14].

generalized [KCY01]. Glassey [Kin94a]. GMRES [CKY99, KCY01, KYC03]. Gradient [Kin81, KOJ88, KOJ89b, KOJ89c, OK87a, OK87b, OJK89]. Greatest [Kin67]. Guide [GKY78, KORY84, KOY89, OJK88].

held [BS84, Lak84]. Honors [KE98].

IFIP [ESY84]. II [BS84]. ImACS [BD92]. Implementation [KGY80, YK95]. Implicit [SK91]. Improve [SK91]. Indefinite [LCK14, CK01]. Integers [Kin67]. interfaces [ESY84]. International [BD92, Lak84]. Introduction [KOJ89a]. issue [AK10]. Iterative [BD92, GKMY78, Kin71a, Kin71b, Kin71c, Kin73b, Kin74b, KY75, KG77, KGMY79, KY79, KGY79, KYG79, KGY80, KY81, KGRY81, KRYG82, KY83, KCO^{+}84, KO85, KOY86b, KCSY88a, KCSY88b, KY91, Kin94b, LCK14, OJK88, OK90, OK94, YK84b, AKK06, CKY99, CK01, Kin71d, Kin73a, Kin74a, KY92, KCY01, KYC03, YJK84, YK85, YK96, BD92, KH90, KE98]. Itpack [KO83, GKMY78, GKY78, KG77, KGMY79, KGY80, KGRY81, KRYG82, KOY82, KY84, KORY85, KY88, KF92a, KF92b, KHY94a, KHY94b, OK87c, YK81, YK84a]. ITPACKV [KORY84, KOY89, KO90, KR91, RK91, RK92].

January [BS84]. Jr. [AK10, KE98, KVW10]. June [Lak84].

L [RSY93]. LAN [KYC00]. LAN/MGMRES [KYC00]. Large [KEG^{+}79, KGY79, KYG79, KGY80, KGRY81, KRYG82, KY83, KCSY88a, KCSY88b, YK90, KY93, LCK14, OK87c, OJK88, YK72, YK81, YOKH85, KH90, PMKL02, YK93]. life [KVW10]. Linear [Kin71a, Kin71b, Kin71c, Kin71d, Kin73b, KEG^{+}79, KGY79, KYG79, KGY80, KGRY81, KRYG82, KY83, KOY84, KOY86a, YK90, KY93, LHKK79a, LHKK79b, LCK14, OK87c, OJK88, YK72, YK81, YOKH85, YHOK85, YK93, AKMK02, AKK06, BD92, CK01, Kin73a, KH90, KYC00, Lak84, PMKL02].

M [AK10, KE98, Kin04, KVW10]. Manual [CK80b, CK85b, KC93b]. Mathematics [BS84, CK80a, CK80b, CK85a, CK85b, CK94a, CK94b, FDKN04, KC91b, KC93b, KC93a, KC96, KC02, KC09]. Matrices [OK88]. Matrix [AEKS89]. Mechanics [BS84]. memory [AK10, RSY93]. Mesh [KY86]. Method [KY71, KY72, Kin72, Kin81, Kin04, OK87a, OK87b, YK69, YK71, CKY99, CK01, KCY01, KYC03]. Methods [Kin71a, Kin71b, Kin71c, Kin71d, Kin73b, Kin74b, Kin74a, KY75, KY79, KGY79, KYG79, KGY80, KGRY81, KRYG82, KCO^{+}84, KO85, KY86, KOY86b, YK90, KY91, KY93, Kin94b, KE98, LCK14, OJK88, YK69, YK72, YK84b, BD92, Kin73a, KH90, KY92, KYC00, KC01, YJK84, YK85, YK93, YK95, YK96]. MGMRES [KYC00]. Model [SK91]. Modifications [LCK14, CKY99]. Modified [KY72, YK71]. Modules [KORY85, ESY84]. Monterey [BS84]. MP [RK91, RK92]. MP4 [KO88b, KO89].

Naval [BS84]. network [PMKL02]. Non [Lak84, CK01]. Non-linear [Lak84]. non-symmetric [CK01]. nonstationary [YK95]. Nonsymmetric [KY81, Kin81, KOJ88, KOJ89b, KOJ89c, OJK89, AKMK02, KYC00]. nonsymmetrizable [YJK84]. Norms [Kin71a, Kin71b, Kin71c, KY71, Kin72, Kin73b, YK69, Kin73a]. NSPCG [KOJ88, KOJ89a, KOJ89b, KOJ89c, OJK88, OJK89]. Numerical [CK80a, CK80b, CK85a, CK85b, CK94a, CK94b, FDKN04, Kin74a, KG77, KC91a, KC91b, KC93b, KC93a, KC96, Kin97, KC02, KC09, OK87a, OK87b, FDKN04, Kin94a].

October [KH90]. Odir [CK01]. Office [BS84]. Oil [KO85, OK87c]. one [RSY93]. Ordering [KY86]. Overrelaxation [KY71, KY72, Kin72, Kin81, Kin04, YK69, YK71]. Overview [KOJ88, KOJ89b, KOJ89c, OJK89, KYC00].

P. [RSY93]. Package [KY75, KGRY81, KRYG82, KOJ88, KOJ89a, KOJ89b, KOJ89c, OJK88, OJK89, YK81, YK84a]. Packages [KEG^{+}79]. Parallel [KO88b, KCSY88a, KCSY88b, KO89, OK87a, OK87b, OK88, SK91, YK95, KO88a, KC01, YK96]. Paralleling [KR91]. Parallelization [KF92a, KF92b, KO90]. Parallelizing [RK91, RK92]. Parameters [KY72, Kin81, YK71]. Partial [KY75, KCO^{+}84, KOY86b, YK84b, YK85]. Past [KY84, KHY94a]. PDE [ESY84, ESY84, KGY79, KYG79, KCSY88a]. PDE-Related [KGY79, KYG79]. PDEs [KCSY88b]. Performance [AKMK02, OK87c, SK91]. Points [KY86]. polynomials [RSY93]. Practice [Lak84]. Prague [FDKN04]. Preconditioned [KOJ88, KOJ89b, KOJ89c, OJK89, AKMK02]. Preface [AK10]. Present [KY84, KHY94a]. problem [BS84].

Problems [KO85, OK87c]. Procedures [Kin67]. Proceedings [BS84, FDKN04, Lak84, BD92, ESY84]. Processor [KO88b, KO89]. project [KY88, KY84].

Recurrences [KY91, KY92]. Reduction [SK91]. refinement [AKK06]. Related [KGY79, YK69, KYG79]. Report [GKMY78, KG77]. Research [BS84, KC01]. Reservoir [KO85, OK87c]. Results [Kin74a]. Review [Kin94a, Kin97, KY88]. Robert [Kin94a]. Routines [KOY82].

Sample [KO85, OK87c]. Scientific [KC91b, KC93b, KC93a, KC96, KC02, KC09]. Second [Kin71d, Kin74b, Kin74a, YK90, KY91, KY93, Kin94b, YK72, KY92, YK93]. Second-Degree [Kin74b, YK90, KY91, KY93, Kin94b, Kin71d, Kin74a, YK72, KY92, YK93]. Semi [Kin71d, Kin74a]. Semi-iterative [Kin71d, Kin74a]. Several [KG77, RSY93]. Simulation [KO85, OK87c]. Söderköping [ESY84]. Software [AEKS89, CK94b, KEG^{+}79, KOJ89a, KF92a, KF92b, KC93a, YK84a, ESY84, KC01]. Solution [CK80b, CK85b, KO85, KORY85, KCSY88a, KCSY88b, YK90, KY93, KC93b, YK72, YK93]. solvers [BS84]. Solving [Kin71c, Kin73b, KY75, KGY79, KYG79, KGY80, KGRY81, KRYG82, KY83, KCO^{+}84, LCK14, OK87c, OJK88, YK84b, YOKH85, YHOK85, AKK06, Kin73a, KYC00, PMKL02, YK85]. Some [KEG^{+}79, KO88b, KO89, LCK14]. Sound [SK91]. Sparse [AEKS89, GKMY78, KGMY79, KEG^{+}79, KGY79, KYG79, KGY80, KGRY81, KRYG82, KY83, KOY84, KOY86a, KCSY88a, KCSY88b, LCK14, OK87c, OJK88, YK81, YOKH85, YHOK85, AKMK02, AKK06, PMKL02]. Special [AK10]. sponsored [BS84]. Standards [AEKS89]. Stationary [YK90, KY91, KY92, KY93, Kin94b, YK72, YK93]. Stewart [Kin97]. Storage [GKMY78, KGMY79]. structure [PMKL02]. Studies [KG77]. Study [KO85]. Subprograms [LHKK79a, LHKK79b]. Successive [KY71, KY72, Kin72, Kin81, Kin04, YK69, YK71]. Supercomputer [KO88b, KO89]. Supercomputers [YK84b, KO83, YK85]. Survey [KY79]. Sweden [ESY84]. Symmetric [GKMY78, KGMY79, KY81, Kin81, LCK14, CK01]. Symposium [BD92, KE98]. Systems [Kin71a, Kin71b, Kin71c, Kin71d, Kin73b, KEG^{+}79, KGY79, KYG79, KGY80, KY81, Kin81, KGRY81, KRYG82, KY83, KOY84, KOY86a, KCSY88a, KCSY88b, YK90, KY93, LCK14, OK87c, OJK88, YK72, YK81, YOKH85, YHOK85, AKMK02, AKK06, CK01, ESY84, Kin73a, KH90, KYC00, PMKL02, YK93].

TC [ESY84]. Texas [KC01, Lak84]. their [RSY93]. Theory [Lak84]. There [OK90, OK94]. times [KVW10]. Topics [RSY93]. Trends [Lak84, Lak84]. Tutorial [KY86]. TX [KH90]. type [YK95, YK96].

underdetermined [PMKL02]. Underwater [SK91]. University [Lak84, KC01]. Usage [LHKK79a, LHKK79b]. Use [KGY79, KYG79, KOY82, KY83, YK84b, YOKH85, YHOK85, YK85]. User [GKY78, KORY84, KOY89, OJK88]. Using [GKMY78, KGMY79, Kin94a, SK91].

variables [RSY93]. Variations [KYC03]. Various [OJK88, AKMK02]. Vector [KOY82, KOY84, KO85, KOY86a, KCSY88a, KCSY88b, OK87c, YOKH85, YHOK85]. vectorization [KO90]. Vectorized [KOY86b]. Version [OJK88]. VIth [Lak84]. volume [RSY93].

W [Kin97]. Working [ESY84]. Workshops [KE98].

X [KO88b, KO89]. X-MP4 [KO88b, KO89].

Y-MP [RK91, RK92]. Years [Kin04]. Young [AK10, KE98, KVW10, Kin04].

References

Ashby:1989:SSM

[AEKS89] S. Ashby, H. Elman, D. Kincaid, and Y. Saad. Standards for sparse matrix software. Summary of session: Standards for Sparse Matrix Software, Copper Mountain Conference on Iterative Methods, April 1990., 1989.

Axelsson:2010:PSI

[AK10] Owe Axelsson and David R. Kincaid. Preface [special issue: dedicated to the memory of David M. Young, Jr.]. Numerical Linear Algebra with Applications, 17 (5):741–742, October 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Al-Kurdi:2006:DIR

[AKK06] Ahmad Al-Kurdi and David R.

Kincaid. LU-decomposition with iterative refinement for solving sparse linear systems. Journal of Computational and Applied Mathematics, 185(2):391– 403, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Al-Kurdi:2002:PVP

[AKMK02] Ahmad Al-Kurdi, R. C. Mittal, and David R. Kincaid. Performance of various preconditioned GCR(k) algorithms applied to sparse nonsymmetric linear systems. Internat. J. Appl. Sci. Comput., 9(1):11–31, 2002. ISSN 1089-0025.

Beauwens:1992:IML

[BD92] R. (Robert) Beauwens and P. (Pieter) De Groen, editors. Iterative methods in linear algebra: proceedings of the ImACS International Symposium on Iterative Methods in Linear Algebra, Brussels, Belgium, 2-4 April, 1991. North-Holland Publishing Co., Amsterdam, The Netherlands, 1992. ISBN 0-444-89248-6. LCCN QA184 .I44 1991.

Birkhoff:1984:EPS

[BS84] Garrett Birkhoff and Arthur L. Schoenstadt, editors. Elliptic problem solvers II: Proceedings of the Elliptic Problem Solvers Conference, sponsored by the Mathematics and Mechanics Branch, Office of Naval Research, held in Monterey, California, January 10-12, 1983. Academic Press, New York, NY, USA, 1984. ISBN 0-12-100560-7. LCCN QA377 .E533 1983.

Cheney:1980:NMC

[CK80a] Ward Cheney and David Kincaid. Numerical Mathematics and Computing. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1980. ISBN 0-8185-0357-2. xiv + 362 pp. LCCN QA297.C426 519.4 79-17230. Contemporary Undergraduate Mathematics Series.

Cheney:1980:SMN

[CK80b] Ward Cheney and David Kincaid. Solution Manual for Numerical Mathematics and Computing. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1980. 212 pp.

Cheney:1985:NMC

[CK85a] Ward Cheney and David Kincaid. Numerical Mathematics and Computing. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, second edition, 1985. ISBN 0-534-04356-9. 562 pp. LCCN QA297.C426 19845.

Cheney:1985:SMN

[CK85b] Ward Cheney and David Kincaid. Solution Manual for Numerical Mathematics and Computing. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1985. ISBN 0-534-04357-7. 292 pp.

Cheney:1994:NMC

[CK94a] Ward Cheney and David Kincaid. Numerical Mathematics and Computing. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, third edition, 1994. ISBN 0-534-20112-1. 578 pp. LCCN QA297.C426 1994.

Cheney:1994:SNM

[CK94b] Ward Cheney and David Kincaid. Software for numerical mathematics and computing, 1994. URL ftp://ftp.

math.utexas.edu/pub/papers/
CNA/cheney-kincaid.

Chronopoulos:2001:OIM

[CK01] Anthony T. Chronopoulos and David Kincaid. On the Odir iterative method for non-symmetric indefinite linear systems. Numerical Linear Algebra with Applications, 8(2):71–82, 2001. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:1999:GMG

[CKY99] Jen-Yuan Chen, David R. Kincaid, and David M. Young. Generalizations and modifications of the GMRES iterative method. Numerical Algorithms, 21(1–4):119–146, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/
content/getfile/5058/20/12/
abstract.htm;   http://ipsapp007.
kluweronline.com/content/getfile/
5058/20/12/fulltext.pdf. Nu-

merical methods for partial differential equations (Marrakech, 1998).

Engquist:1984:PSM

[ESY84] Björn Engquist, Tom Smedsaas, and N. N. (Nikolaĭ Nikolaevich) Yanenko, editors. PDE software: modules, interfaces, and systems: proceedings of the IFIP TC 2 Working Conference on PDE Software–Modules, Interfaces, and Systems, Söderköping, Sweden, 22–26 August, 1983. North-Holland Publishing Co., Amsterdam, The Netherlands, 1984. ISBN 0-444-87620-0. LCCN QA377 .I44 1983.

Feistauer:2004:NMA

[FDKN04] M. (Miloslav) Feistauer, Vít Dolejší, Petr Knobloch, and Karel Najzar, editors. Numerical Mathematics and Advanced Applications: Proceedings of ENU-MATH 2003 the 5th European Conference on Numerical Mathematics and Advanced Applications Prague, August 2003. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-642-62288-7, 3-540-21460-7, 3-642-18775-7 (e-book). LCCN QA297 .E89 2003. URL http:/

/link.springer.com/10.1007/
978-3-642-18775-9.

Grimes:1978:IRA

[GKMY78] Roger G. Grimes, David R. Kincaid, William I. MacGregor, and David M. Young. ITPACK report: Adaptive iterative algorithms using symmetric sparse storage. Report CNA-139, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, August 1978.

Grimes:1978:IUG

[GKY78] Roger G. Grimes, David R. Kincaid, and David M. Young. IT-PACK 2.0 user’s guide. Report CNA-150, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, August 1978.

Kincaid:1991:NA

[KC91a] David Kincaid and Ward Cheney.

Numerical analysis. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1991. ISBN 0-534-13014-3. x + 690 pp. Mathematics of scientific computing.

Kincaid:1991:NAM

[KC91b] David (David Ronald) Kincaid and E. W. Cheney. Numerical Analysis: Mathematics of Scientific Computing. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1991. ISBN 0-534-13014-3. viii + 690 pp. LCCN QA297 .K563 1990. URL http:/

                                        /www.zentralblatt-math.org/
                                        zmath/en/search/?an=0745.65001.

Kincaid:1993:SNA

[KC93a] David Kincaid and Ward Cheney. Software for numerical analysis: Mathematics of scientific computing, 1993. URL ftp://ftp.

                                        math.utexas.edu/pub/papers/
                                        CNA/kincaid-cheney.

Kincaid:1993:SMN

[KC93b] David Kincaid and Ward Cheney. Solution Manual for Numerical Analysis: Mathematics of Scientific Computing. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1993. 219 pp.

Kincaid:1996:NAM

[KC96] David (David Ronald) Kincaid and E. W. Cheney. Numerical analysis: mathematics of scientific computing. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, second edition, 1996. ISBN 0-534-33892-5. xii + 804 pp. LCCN QA297 .K563 1996. Mathematics of scientific computing.

Kincaid:2001:RPM

[KC01] David R. Kincaid and Graham F.

Carey. Research on parallel methods, algorithms, and software at the University of Texas at Austin. Mat. Model., 13(3):11–21, 2001. ISSN 0234-0879.

Kincaid:2002:NAM

[KC02] David (David Ronald) Kincaid and E. W. (Elliott Ward) Cheney. Numerical analysis: mathematics of scientific computing. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, third edition, 2002. ISBN 0-534-38905-8. xiv + 788 pp. LCCN QA297 .K563 2002.

Kincaid:2009:NAM

[KC09] David (David Ronald) Kincaid and E. W. (Elliott Ward) Cheney. Numerical analysis: mathematics of scientific computing, volume 2 of The Sally series; Pure and applied undergraduate texts. American Mathematical Society, Providence, RI, USA, third edition, 2009. ISBN 0-8218-4788-0. xiv + 788 pp. LCCN QA297 .K563 2009. URL http://www.

loc.gov/catdir/toc/fy0906/
2008047389.html.

Kincaid:1984:CFE

[KCO^{+}84] David R. Kincaid, Graham F.

Carey, Thomas C. Oppe, Kamy Sepehrnoori, and David M. Young. Combining finite element and iterative methods for solving partial differential equations on advanced computer architectures. In R. Vichnevetsky and R. S. Stepleman, editors, Advances in Computer Methods for Partial Differential Equations, volume V, pages 375–378. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1984.

Kincaid:1988:VPIa

[KCSY88a] David R. Kincaid, Graham F. Carey, Kamy Sepehrnoori, and David M. Young. Vector and parallel iterative solution of large sparse systems for PDE’s. In Science and Engineering on Cray Supercomputers, pages 25–44. Cray Research, Inc., Minneapolis, MN, 1988.

Kincaid:1988:VPIb

[KCSY88b] David R. Kincaid, Graham F. Carey, Kamy Sepehrnoori, and David M. Young. Vector and parallel iterative solution of large sparse systems for PDEs. Report CNA-222, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, August 1988.

Kincaid:2001:GGI

[KCY01] David R. Kincaid, Jen-Yuan Chen, and David M. Young. A generalized GMRES iterative method. In Numerical analysis and its applications (Rousse, 2000), volume 1988 of Lecture Notes in Computer Science, pages 475–481. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc.,

Kincaid:1998:CWI

[KE98] David R. Kincaid and Anne C.

Elster. Conferences & workshops: Iterative Methods Symposium Honors David M. Young, Jr. IEEE Computational Science & Engineering, 5(4):12–15, October/December 1998. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1998/pdf/c4011.
pdf.

Kincaid:1979:SCS

[KEG^{+}79] D. Kincaid, S. Eisenstat, A. George, R. Grimes, and A. Sherman. Some comparisons of software packages for large sparse linear systems. In R. Vichnevetsky and R. S. Stepleman, editors, Advances in Computer Methods for Partial Differential Equations, volume III, pages 98–106. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1979.

Kincaid:1992:ISPa

[KF92a] David R. Kincaid and Matthew D.

Fassiotto. ITPACK software and parallelization. In D. Knight R. Vichnevetsky and G. Richter, editors, Advances in Computer Methods for Partial Differential Equations, volume VII, pages 425–430. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1992.

Kincaid:1992:ISPb

[KF92b] David R. Kincaid and Matthew D. Fassiotto. ITPACK software and parallelization. Report CNA-255, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, June 1992.

Kincaid:1977:IRN

[KG77] David R. Kincaid and Roger G. Grimes. ITPACK report: Numerical studies of several adaptive iterative algorithms. Report CNA-126, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, August 1977.

Kincaid:1979:IAI

[KGMY79] David R. Kincaid, Roger G. Grimes, William I. MacGregor, and David M. Young. ITPACK: Adaptive iterative algorithms using symmetric sparse storage. In Symposium on Reservoir Simulation, volume SPE 7687, pages 151–170. Society of Petroleum Engineers of AIME, Dallas, TX, 1979.

Kincaid:1981:IFP

[KGRY81] David R. Kincaid, Roger G. Grimes, John R. Respess, and David M. Young. ITPACK 2B: A Fortran package for solving large sparse linear systems by adaptive accelerated iterative methods. Report CNA-173, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, September 1981. (Also, Report CCSN–44, Computation Center, University of Texas at Austin.).

Kincaid:1979:UIMa

[KGY79] David R. Kincaid, Roger G.

Grimes, and David M. Young. The use of iterative methods for solving large sparse PDE-related linear systems. Mathematics and Computers in Simulation, XXI: 368–375, 1979. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Kincaid:1980:IFI

[KGY80] David R. Kincaid, Roger G.

Grimes, and David M. Young. IT-PACK 2A: A Fortran implementation of adaptive accelerated iterative methods for solving large sparse linear systems. Report CNA-164, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, October 1980.

Kincaid:1990:IML

[KH90] David R. Kincaid and Linda J.

Hayes, editors. Iterative methods for large linear systems (Austin, TX, October 19–21, 1988). Academic Press, New York, NY, USA, 1990. ISBN 0-12-407475-8. LCCN QA432 .I84 1990. Papers from a conference held Oct. 19– 21, 1988, at the Center for Numerical Analysis of the University of Texas at Austin. Banquet in honour of David M. Young’s 65th birthday.

Kincaid:1994:IPP

[KHY94a] David R. Kincaid, Linda J.

Hayes, and David M. Young. IT-PACK: Past, present, and future. In Colorado Conference on Iterative Methods, volume 1, page ?? University of Colorado & Front Range Scientific Computations, Inc., Boulder, CO, 1994.

Kincaid:1994:ITN

[KHY94b] David R. Kincaid, Linda J. Hayes, and David M. Young. IT-PACK: Then and now. In IMACS 13th World Congress on Computational and Applied Mathematics. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1994. To appear.

Kincaid:1967:APC

[Kin67] David R. Kincaid. Algorithmic procedures for computing the greatest common divisor of n integers. Master’s thesis, University of Texas at Austin, Austin, TX, USA, May 1967.

Kincaid:1971:ACNa

[Kin71a] David R. Kincaid. An analysis of a class of norms of iterative methods for systems of linear equations. Report CNA-18, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, May 1971. Ph.D. thesis.

Kincaid:1971:ACNb

[Kin71b] David R. Kincaid. An Analysis of a Class of Norms of Iterative Methods for Systems of Linear Equations. PhD thesis, University of Texas at Austin, May 1971. Also, Report CNA-18, Center for Numerical Analysis.

Kincaid:1971:CNI

[Kin71c] David R. Kincaid. A class of norms of iterative methods for solving systems of linear equations. Report CNA-24, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, June 1971.

Kincaid:1971:ESI

[Kin71d] David R. Kincaid. Extrapolated, semi-iterative and seconddegree methods for systems of linear equations. Report CNA-23, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, June 1971.

Kincaid:1972:NSO

[Kin72]  David R. Kincaid.  Norms

of the successive overrelaxation method. Mathematics of Computation, 26(118):345–357, 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kincaid:1972:CNI

[Kin73a] David R. Kincaid. A class of norms of iterative methods for solving systems of linear equations. Numerische Mathematik, 20:392–408, 1972/1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kincaid:1973:CNI

[Kin73b] David R. Kincaid. A class of norms of iterative methods for solving systems of linear equations. Numerische Mathematik, 20:392–408, 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kincaid:1974:NRA

[Kin74a] David R. Kincaid. Numerical results of the application of complex second-degree and semiiterative methods. Report CNA-90, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, September 1974.

Kincaid:1974:CSD

[Kin74b] David R. Kincaid. On complex second-degree iterative methods. SIAM Journal on Numerical Analysis, 11(2):211–218, April 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kincaid:1981:APS

[Kin81] David R. Kincaid. Acceleration parameters for a symmetric successive overrelaxation conjugate gradient method for nonsymmetric systems. In R. Vichnevetsky and R. S. Stepleman, editors, Advances in Computer Methods for Partial Differential Equations, volume IV, pages 294–299. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1981.

Kincaid:1994:BRB

[Kin94a] David R. Kincaid. Book review: Numerical Computation Using C (Robert Glassey). SIAM Review, 36(4):672–673, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kincaid:1994:SSD

[Kin94b] David R. Kincaid. Stationary second-degree iterative methods. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2):227–237, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). A Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Kincaid:1997:BRA

[Kin97] David R. Kincaid. Book review:

Afternotes on Numerical Analysis, by G. W. Stewart. SIAM Review, 39(1):153, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kincaid:2004:CFY

[Kin04] David R. Kincaid. Celebrating fifty years of David M. Young’s successive overrelaxation method. In Feistauer et al. [FDKN04], pages 549–558. ISBN 3-642-62288-7, 3-540-21460-7, 3-642-18775-7 (e-book). LCCN QA297 .E89 2003. URL http:/

/link.springer.com/10.1007/
978-3-642-18775-9.

Kincaid:1983:IS

[KO83] David R. Kincaid and Thomas C.

Oppe. Itpack on supercomputers. Lecture Notes in Mathematics, 1005:151–161, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0112532/.

Kincaid:1985:CSI

[KO85] David R. Kincaid and Thomas C. Oppe. A comparison study of iterative solution methods for sample oil reservoir simulation problems on vector computers. Report CNA-200, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, August 1985.

Kincaid:1988:PAG

[KO88a] David R. Kincaid and Thomas C. Oppe. A parallel algorithm for the general LU factorization. Communications in Applied Numerical Methods, 4(3): 349–359, 1988. CODEN CAN-MER. ISSN 0748-8025 (print), 1555-2047 (electronic).

Kincaid:1988:SPA

[KO88b] David R. Kincaid and Thomas C. Oppe. Some parallel algorithms on the four processor Cray X-MP4 supercomputer. Report CNA-220, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, May 1988.

Kincaid:1989:SPA

[KO89] David R. Kincaid and Thomas C. Oppe. Some parallel algorithms on the four processor Cray X-MP4 supercomputer. In Graham F. Carey, editor, Parallel Supercomputing: Methods, Algorithms and Applications, pages 121–134. Wiley, New York, NY, USA, 1989. ISBN 0-471-92436-9.

Kincaid:1990:RVP

[KO90] David R. Kincaid and Thomas C.

Oppe. Recent vectorization and parallelization of ITPACKV. Lecture Notes in Mathematics, 1457: 58–78, 1990. CODEN LN-MAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0090902/.

Kincaid:1988:ONN

[KOJ88] David R. Kincaid, Thomas C.

Oppe, and Wayne D. Joubert. An overview of NSPCG: A nonsymmetric preconditioned conjugate gradient package. Report CNA-228, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, October 1988.

Kincaid:1989:INS

[KOJ89a] David R. Kincaid, Thomas C.

Oppe, and Wayne D. Joubert. An introduction to the NSPCG software package. International J. Numer. Methods Engineering, 27 (3):589–608, 1989.

Kincaid:1989:ONNa

[KOJ89b] David R. Kincaid, Thomas C.

Oppe, and Wayne D. Joubert. An overview of NSPCG: A nonsymmetric preconditioned conjugate gradient package. In Daniel L. Boley, Donald G. Truhlar, Youcef Saad, Robert E. Wyatt, and Lee A. Collins, editors, Practical Iterative Methods for Large Scale Computations, pages 283–293. North-Holland Publishing Co., Amsterdam, The Netherlands, 1989. ISBN 0-444-88023-2. LCCN QA297.8.M56 1988. (Reprinted from Computer Physics Communications, 53(3):283–293, 1989.).

Kincaid:1989:ONNb

[KOJ89c] David R. Kincaid, Thomas C. Oppe, and Wayne D. Joubert. An overview of NSPCG: a nonsymmetric preconditioned conjugate gradient package. Computer Physics Communications, 53(1–3):283–293, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Kincaid:1984:IUG

[KORY84] David R. Kincaid, Thomas C. Oppe, John R. Respess, and David M. Young. ITPACKV 2C user’s guide. Report CNA-191, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, February 1984.

Kincaid:1985:ISM

[KORY85] David R. Kincaid, Thomas C. Oppe, John R. Respess, and David M. Young. ITPACK solution modules. In John R. Rice and Ronald F. Boisvert, editors, Solving Elliptic Problems Using ELLPACK, pages 121– 146. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-90910-9. LCCN QA377.R53 1984.

Kincaid:1982:AIR

[KOY82] David R. Kincaid, Thomas C. Oppe, and David M. Young.

Adapting ITPACK routines for use on a vector computer. In Proceedings: Symposium on CY-BER 205 Applications, page ?? Colorado State University, Fort Collins, CO, August 1982.

Kincaid:1984:VCS

[KOY84] David R. Kincaid, Thomas C.

Oppe, and David M. Young. Vector computations for sparse linear systems. Report CNA-189, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, February 1984.

Kincaid:1986:VCS

[KOY86a] David R. Kincaid, Thomas C.

Oppe, and David M. Young. Vector computations for sparse linear systems. SIAM Journal on Algebraic and Discrete Methods, 7(1):99–112, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kincaid:1986:VIM

[KOY86b] David R. Kincaid, Thomas C.

Oppe, and David M. Young. Vectorized iterative methods for partial differential equations. Communications in Applied Numerical Methods, 2(3):289–296, 1986. CODEN CANMER. ISSN 0748-8025 (print), 1555-2047 (electronic).

Kincaid:1989:IUG

[KOY89] David R. Kincaid, Thomas C.

Oppe, and David M. Young. IT-PACKV 2D user’s guide. Report CNA-232, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, May 1989.

Kincaid:1991:PI

[KR91] David R. Kincaid and Malathi Ramdas. Paralleling ITPACKV 2D. In J. J. H. Miller and R. Vichnevetsky, editors, IMACS ’91: Proceedings of the 13th IMACS World Congress on Computational and Applied Mathematics, volume 2, pages 686–687. Criterion Press, Dublin, Ireland, 1991.

Kincaid:1982:AIF

[KRYG82] David R. Kincaid, John R. Respess, David M. Young, and Roger G. Grimes. Algorithm 586: ITPACK 2C: A FORTRAN package for solving large sparse linear systems by adaptive accelerated iterative methods. ACM Transactions on Mathematical Software, 8(3):302–322, 1982. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kincaid:2010:LTD

[KVW10] David R. Kincaid, Richard S. Varga, and Charles H. Warlick. The life and times of Dr. David M. Young, Jr. Numerical Linear Algebra with Applications, 17 (5):743–757, October 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kincaid:1971:NSO

[KY71] David R. Kincaid and David M. Young. Norms of the successive overrelaxation method. Report CNA-26, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, July 1971.

Kincaid:1972:MSO

[KY72] David R. Kincaid and David M.

Young. The modified successive overrelaxation method with fixed parameters. Mathematics of Computation, 26(119):705– 717, 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kincaid:1975:DCP

[KY75] David R. Kincaid and David M.

Young. The development of a computer package for solving a class of partial differential equations by iterative methods. Annales de l’Association Internationale pour le Calcul Analogique, 3:186–191, 1975. CO-DEN AAICAE. ISSN 0020-594X.

Kincaid:1979:SIM

[KY79] David R. Kincaid and David M.

Young. Survey of iterative methods. In Jack Belzer, Albert G. Holzman, and Allen Kent, editors, Encyclopedia of Computer Sciences and Technology, volume 13, pages 354–391. Marcel Dekker, New York, NY, USA, 1979. ISBN 0-8247-2263-9.

Kincaid:1981:AIA

[KY81] David R. Kincaid and David M.

Young. Adapting iterative algorithms developed for symmetric systems to nonsymmetric systems. In Martin Schultz, editor, Elliptic Problem Solvers, pages 353–359. Academic Press, New York, NY, USA, 1981. ISBN 0-12-632620-7. LCCN QA377.E53.

Kincaid:1983:AIA

[KY83] David R. Kincaid and David M. Young. Adapting iterative algorithms for solving large sparse linear systems for efficient use on the CDC CYBER 205. In CYBER 200 Applications Seminar, pages 147–160. NASA Scientific and Technical Information Research, Washington, D.C., 1983. NASA Conference Publication 2295.

Kincaid:1984:IPPa

[KY84] David R. Kincaid and David M. Young. The ITPACK Project: Past, present, and future. In Birkhoff and Schoenstadt [BS84], pages 53–63. ISBN 0-12-100560-7. LCCN QA377 .E533 1983.

Kincaid:1986:TFD

[KY86] David R. Kincaid and David M. Young. A tutorial on finite difference methods and ordering of mesh points. In H. S. Stone and S. Winkler, editors, Fall Joint Computer Conference Proceedings, pages 556–559. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. ISBN 0-8186-0743-2.

Kincaid:1988:BRI

[KY88] David R. Kincaid and David M. Young. A brief review of the ITPACK project. Journal of Computational and Applied Mathematics, 24(1–2):121– 127, November 1988. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0377042788903470.

Kincaid:1991:SSD

[KY91] David R. Kincaid and David M.

Young. Stationary second-degree iterative methods and recurrences. Report CNA-250, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, February 1991.

Kincaid:1992:SSD

[KY92] David R. Kincaid and David M.

Young. Stationary seconddegree iterative methods and recurrences. In Beauwens and De Groen [BD92], pages 27– 47. ISBN 0-444-89248-6. LCCN QA184 .I44 1991.

Kincaid:1993:LSS

[KY93] David R. Kincaid and David M.

Young. Linear stationary second-degree methods for solution of large linear systems. In Rassias et al. [RSY93], pages 609–629. ISBN 981-02-0614-3. LCCN QA161.P59 T66 1993. URL http://

www.zentralblatt-math.org/
zmath/en/search/?an=0849.00029.

Kincaid:2000:OML

[KYC00] David R. Kincaid, David M.

Young, and Jen-Yuan Chen. An overview of MGMRES and LAN/ MGMRES methods for solving nonsymmetric linear systems. Taiwanese J. Math., 4(3):385– 396, 2000. ISSN 1027-5487.

Kincaid:2003:VGI

[KYC03] David R. Kincaid, David M. Young, and Jen-Yuan Chen. Variations of the GMRES iterative method. Applied Numerical Mathematics: Transactions of IMACS, 45(1):3–10, 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). 5th IMACS Conference on Iterative Methods in Scientific Computing (Heraklion, 2001).

Kincaid:1979:UIMb

[KYG79] David R. Kincaid, David M. Young, and Roger G. Grimes. The use of iterative methods for solving large sparse PDE-related linear systems. In R. Vichnevetsky and R. S. Stepleman, editors, Advances in Computer Methods for Partial Differential Equations, volume III, pages 29– 32. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1979.

Lakshmikantham:1984:TTP

[Lak84] V. Lakshmikantham, editor. Trends in the Theory and Practice of Non-linear Analysis: Proceedings of the VIth International Conference on Trends in the Theory and Practice of Non-Linear Analysis held at the University of Texas at Arlington, June 18–22, 1984. Elsevier Science Publishers (North-Holland), New York, NY, USA, 1984. ISBN 0-444-87704-5. LCCN QA299.6 .I57 1984.

Li:2014:SGM

[LCK14] Yu-Chien Li, Jen-Yuan Chen, and David R. Kincaid. Some generalizations and modifications of iterative methods for solving large sparse symmetric indefinite linear systems. Abstract and Applied Analysis, page 11, 2014. ISSN 1085-3375 (print), 1687-0409 (electronic). Article ID 237808.

Lawson:1979:ABL

[LHKK79a] C. L. Lawson, R. J. Hanson, D. R.

Kincaid, and F. T. Krogh. Algorithm 539: Basic Linear Algebra Subprograms for Fortran usage [F1]. ACM Transactions on Mathematical Software, 5(3): 324–325, September 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lawson:1979:BLA

[LHKK79b] C. L. Lawson, R. J. Hanson, D. R.

Kincaid, and F. T. Krogh. Basic Linear Algebra Subprograms for Fortran usage. ACM Transactions on Mathematical Software, 5(3):308–323, September 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Oppe:1988:NUG

[OJK88] Thomas C. Oppe, Wayne D.

Joubert, and David R. Kincaid. NSPCG user’s guide, version 1.0: A package for solving large sparse linear systems by various iterative methods. Report CNA-216, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, April 1988.

Oppe:1989:ONN

[OJK89] Thomas C. Oppe, Wayne D. Joubert, and David R. Kincaid. An overview of NSPCG: a nonsymmetric preconditioned conjugate gradient package. Computer Physics Communications, 53(1–3):283–293, May 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0010465589901665.

Oppe:1987:NEPa

[OK87a] T. Oppe and D. Kincaid. Numerical experiments with a parallel conjugate gradient method. Technical report CNA-208, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, 1987.

Oppe:1987:NEPb

[OK87b] Thomas C. Oppe and David R. Kincaid. Numerical experiments with a parallel conjugate gradient method. In R. Vichnevetsky and R. S. Stepleman, editors, Advances in Computer Methods for Partial Differential Equations, volume VI, pages 369–374. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1987.

Oppe:1987:PIV

[OK87c] Thomas C. Oppe and David R. Kincaid. The performance of IT-PACK on vector computers for solving large sparse linear systems arising in sample oil reservoir simulation problems. Communications in Applied Numerical Methods, 3(1):23–29, 1987. CODEN CANMER. ISSN 0748-8025 (print), 1555-2047 (electronic).

Oppe:1988:PFA

[OK88] Thomas C. Oppe and David R.

Kincaid. Parallel LU-factorization algorithms for dense matrices. In E. N. Houstis, T. S. Papatheodorou, and C. D. Polychronopoulos, editors, Supercomputing (Athens, 1987), volume 297 of Lecture Notes in Computer Science, pages 576–594. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-18991-2.

Oppe:1990:TIB

[OK90] Thomas C. Oppe and David R.

Kincaid. Are there iterative BLAS? Report CNA-240, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, February 1990.

Oppe:1994:TIB

[OK94] Thomas C. Oppe and David R.

Kincaid. Are there iterative BLAS? International Journal on Scientific Computing & Modeling, 1:??, 1994. To appear.

Pilipchuk:2002:ASL

[PMKL02] Ludmila A. Pilipchuk, Yulia V.

Malakhouskaya, David R. Kincaid, and Minghorng Lai. Algorithms of solving large sparse underdetermined linear systems with embedded network structure. East-West J. Math., 4(2): 191–201, 2002. ISSN 1513-489X.

Ramdas:1991:PIC

[RK91] Malathi Ramdas and David R. Kincaid. Parallelizing ITPACKV 2D for the Cray Y-MP. Report CNA-249, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, February 1991.

Ramdas:1992:PIC

[RK92] Malathi Ramdas and David R. Kincaid. Parallelizing ITPACKV 2D for the Cray Y-MP. In Beauwens and De Groen [BD92], pages 323–337. ISBN 0-444-89248-6. LCCN QA184 .I44 1991.

Rassias:1993:TPO

[RSY93] Th. M. Rassias, H. M. Srivastava, and A. Yanushauskas, editors. Topics in polynomials of one and several variables and their applications: volume dedicated to the memory of P. L. Chebyshev (1821–1894). World Scientific Publishing Co., Singapore, Philadelphia, River Edge, NJ, 1993. ISBN 981-02-0614-3. ix + 638 pp. LCCN QA161.P59 T66 1993. URL http://

                                        www.zentralblatt-math.org/
                                        zmath/en/search/?an=0849.00029.

Santiago:1991:UCR

[SK91] Marilyn Santiago and David R. Kincaid. Using cyclic reduction on a parallel computer to improve the performance of an underwater sound implicit finite difference model. Computers in Mathematical Applications, 21(5):83– 94, 1991.

Young:1985:UVCb

[YHOK85] David M. Young, Linda J. Hayes, Thomas C. Oppe, and David R. Kincaid. On the use of vector computers for solving sparse linear systems. In Proceedings of the Conference Vector and Parallel Processors for Scientific Computation, page ?? Accademia dei Lincei, Roma, Italy, May 1985.

Young:1984:ANI

[YJK84] David M. Young, Kang C. Jea, and David R. Kincaid. Accelerating nonsymmetrizable iterative methods. In Birkhoff and Schoenstadt [BS84], pages 323– 342. ISBN 0-12-100560-7. LCCN QA377 .E533 1983.

Young:1969:NSO

[YK69] David M. Young and David R.

Kincaid. Norms of the successive overrelaxation method and related methods. Report TNN– 94, Computation Center, University of Texas at Austin, Austin, TX, USA, September 1969.

Young:1971:MSO

[YK71] David M. Young and David R.

Kincaid. The modified successive overrelaxation method with fixed parameters. Report CNA-33, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, October 1971.

Young:1972:LSS

[YK72] David M. Young and David R. Kincaid. Linear stationary second-degree methods for the solution of large linear systems. Report CNA-52, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, October 1972.

Young:1981:IPL

[YK81] David M. Young and David R. Kincaid. The ITPACK package for large sparse linear systems. In Martin Schultz, editor, Elliptic Problem Solvers, pages 163–185. Academic Press, New York, NY, USA, 1981. ISBN 0-12-632620-7. LCCN QA377.E53.

Young:1984:ISP

[YK84a] David M. Young and David R. Kincaid. The ITPACK software package. In Engquist et al. [ESY84], pages 193–206. ISBN 0-444-87620-0. LCCN QA377 .I44 1983. With a discussion.

Young:1984:UIM

[YK84b] David M. Young and David R. Kincaid. On the use of iterative methods with supercomputers for solving partial differential equations. In Lakshmikantham [Lak84], pages 455–466. ISBN 0-444-87704-5. LCCN QA299.6 .I57 1984.

Young:1985:UIM

[YK85] David M. Young and David R. Kincaid. On the use of iterative methods with supercomputers for solving partial differential equations. In Trends in the theory and practice of nonlinear analysis (Arlington, Tex., 1984), volume 110 of North-Holland Math. Stud., pages 455–466. North-Holland Publishing Co., Amsterdam, The Netherlands, 1985.

Kincaid:1990:LSS

[YK90] David M. Young and David R.

Kincaid. Linear stationary second-degree methods for solution of large linear systems. Report CNA-244, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, July 1990.

Young:1993:LSS

[YK93] David M. Young and David R.

Kincaid. Linear stationary second-degree methods for the solution of large linear systems. In Rassias et al. [RSY93], pages 609–629. ISBN 981-02-0614-3. LCCN QA161.P59 T66 1993. URL http://

www.zentralblatt-math.org/
zmath/en/search/?an=0849.00029.

Young:1995:PIC

[YK95] David M. Young and David R.

Kincaid. Parallel implementation of a class of nonstationary alternating-type methods. In Proceedings of the Third International Colloquium on Numerical Analysis (Plovdiv, 1994), pages 219–222. VSP, Utrecht, 1995.

Young:1996:NCP

[YK96] David M. Young and David R.

Kincaid. A new class of parallel alternating-type iterative methods. Journal of Computational and Applied Mathematics, 74(1–2):331–344, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). TICAM Symposium (Austin, TX, 1995).

Young:1985:UVCa

[YOKH85] D. Young, T. Oppe, D. Kincaid, and L. Hayes. On the use of vector computers for solving large sparse linear systems. Technical report CNA-199, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, 1985.